OpenStudy (vikstar2.0):

Year 1960 1975 1990 2005 2020 (est.) Population P(t) 200,000 240,000 288,000 345,600 414,720 According to the data, what would the population be in 2010? [Note: The function is P(t) = (200)(1.01223)t . A) 367,274 B) 367,987 Eliminate C) 368,254 D) 368,640

OpenStudy (vikstar2.0):

a medal and a fan to who ever can answer it

OpenStudy (vikstar2.0):

@ganeshie8

OpenStudy (vikstar2.0):

@jabez177

OpenStudy (vikstar2.0):

@TheSmartOne

OpenStudy (vikstar2.0):

@ILovePuppiesLol

OpenStudy (mathmale):

Please note that one has to look twice (or more) at your " P(t) = (200)(1.01223)t" to realize that you most likely mean the exponential function\[ P(t) = (200)(1.01223)^t\]

OpenStudy (mathmale):

Also, even if the original problem were given as P(t) = (200)(1.01223)^t, you may have to replace that '200' with 200,000. Note that you represent the year 1960 with t=0, the year 1975 with t=15, and so on. Now this makes more sense: t=0 at the very start. Thus, 15 years later, in 1975, the population would be P(t) = (20,000)(1.01223)^15, or 240,000, which agrees with your table. How would you now predict the population in corresponding to the year 2010? Your main challenge here is to identify the correct exponent to use on the base (1.01223).

OpenStudy (wolf1728):

Hmm, it seems that Vickstar is offline Anyway, as each year passes you add another year to the exponent, so for 2005, the exponent is 45; for 2006 it is 46; for 2007 it is 47 From that information, what is the exponent for 2010?